if Z1 = - 3 + 5 i and Z 2 = - 5 - 3 i and z is a complex number lying on the line segment joining Z1 and Z2 that argument that can be

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
z1 – z2 = – 8 + 8i
 
hence arg z = tan-1 8/-8 = tan-1 (-1) = – \pi/4
 
hope it helps
 
Regards
Arun (askIITians forum expert)